OpenStudy (anonymous):

whats the midpoint between the points P(10, -5) and Q(8, -7).

OpenStudy (aonz):

use midpoint formula http://geometry.jdeer.com/Images/c1/midpoint%20formula.gif

hero (hero):

@AonZ, what's wrong with just typing the midpoint formula: \[m = \left(\frac{x_1 + x_2}{2}, \frac{y_1 + y_2}{2}\right)\]

OpenStudy (aonz):

i dont know latex or anything

OpenStudy (aonz):

it will look horribly messy

hero (hero):

It will let you preview before you post.

hero (hero):

For a fraction you type `\[\frac{a}{b}\]`

hero (hero):

For Parentheses you type `\[\left(\right)\]`

hero (hero):

Try it @celecity

OpenStudy (anonymous):

\[\frac{10+8}{2}\] Woo hoo fractions

hero (hero):

What took you so long?

OpenStudy (anonymous):

I was experimenting with the parentheses and how to combine the two

hero (hero):

\[\frac{\left(10+8\right)}{2}\]

hero (hero):

To see the code, Right Click (on the expression)--> Show Math As --> Tex Commands

hero (hero):

You learn by looking at other people's code.